# Incorporating Magnetic Field Characteristics into EUV-Based Automated Segmentation of Coronal Holes

### v1

@@ Abstract
Coronal holes (CH) are magnetically open regions that allow hot coronal plasma to escape from the Sun and form the high-speed solar wind. This wind can interact with Earth’s magnetic field. For this reason, developing an accurate understanding of CH regions is vital for understanding space weather and its effects on Earth. The process of identifying CH regions typically relies on extreme ultraviolet (EUV) imagery, leveraging the fact that CHs appear dark at these wavelengths. Accurate identification of CHs in EUV, however, can be difficult due to a variety of factors, including stray light from nearby regions, limb brightening, and the presence of filaments (which also appear dark, but are not sources of solar wind). In order to overcome these issues, this work incorporates photospheric magnetic field data into a classical EUV-based segmentation algorithm based on the active contours without edges (ACWE) segmentation method. In this work magnetic field data are incorporated directly into the segmentation process, serving both as a method for removing non-CH regions in advance, and as a method to constrain evolution of the boundary which identifies the CH boundary. This reduces the presence of filaments while simultaneously allowing the segmentation to include CH regions that may be difficult to identify due to the aforementioned inconsistent intensities.
@@ 1 Introduction
Due to the lower density and temperature of CH regions, which is a byproduct of the open magnetic field lines, CHs can be detected in extreme ultraviolet (EUV) imagery, where they appear as dark regions (Munro and Withbroe, 1972). This has resulted in a large number of detection schemes which rely on EUV imagery, including those developed by Krista and Gallagher (2009); Verbeeck et al. (2014); Lowder et al. (2014); Caplan et al. (2016); Boucheron et al. (2016); Garton et al. (2018); Hamada et al. (2018); Grajeda et al. (2023). Each of these methods is developed to address difficulties with the segmentation process, including overcoming inconsistent intensities caused by stray light from nearby regions, limb brightening, and instrument effects (Verbeeck et al., 2014; Caplan et al., 2016). In addition to these difficulties, the presence of filaments, which also appear dark in EUV but are not sources of solar wind, can result in the misclassification of regions when using only EUV data (Krista and Gallagher, 2009; Reiss et al., 2023).
@@
The open magnetic field lines present in CHs emerge from highly unipolar regions unlike magnetic fields in filaments which are bipolar in nature. This has resulted in several methods that validate CH regions after detection through the use of magnetic field data, removing non-CH regions as a post-hoc process in the pipeline (Scholl and Habbal, 2008; Krista and Gallagher, 2009; Lowder et al., 2014; Garton et al., 2018; Hamada et al., 2018; Landeros et al., 2024). With that noted, direct application of magnetic field data during the initial detection and segmentation of CH regions remains fairly limited, appearing predominately in Jarolim et al. (2021) where it serves as an additional channel alongside seven EUV channels input to a network based on the U-Net architecture of Ronneberger et al. (2015). The work presented here proposes the incorporation of magnetic field data into the classical segmentation method first employed by Boucheron et al. (2016) and extended in Grajeda et al. (2023) which is based on the Active Contours Without Edges (ACWE) segmentation process of Chan and Vese (2001). The ACWE-based method has already been shown to effectively overcome the inconsistent intensity present in CH observations for the majority of cases and is shown to be robust across short timeframes where CH evolution is expected to be minimal (Grajeda et al., 2023). However, like all other EUV-based methods, it can misidentify filament regions (Grajeda et al., 2023; Reiss et al., 2023). Through the addition of magnetic field data in both seeding and iterating of the ACWE process, the method proposed here reduces filament contamination, first by eliminating many filaments outright and second by reducing the overall area of the filaments that cannot be outright removed. Additionally, incorporation of magnetic field in the seeding and iterating of ACWE better constrains the evolution to avoid including too much quiet Sun (QS) as CH.
@@
The rest of this work is organized as follows: Section 2 provides an introduction to the ACWE segmentation, outlines how this method was adapted for CH detection, and introduces the dataset that will be used to validate this work. Section 3 outlines the process of incorporating magnetic field data into ACWE evolution and explores the resulting effects. Section 4 outlines the process of incorporating magnetic field data into the initial seeding to help remove filament contamination outright. The results from Sections 3 and 4 are used to develop a final pipeline, which is discussed and validated in Section 5. Conclusions and discussion of future work is provided in Section 6.
@@ 2.1 Active Contours Without Edges (ACWE)
The detection method proposed here is an extension of the ACWE algorithm developed by Chan and Vese (2001) and first applied to CH detection in Boucheron et al. (2016). At its most basic, ACWE is a two-step process designed to separate an image into foreground (in our case the CHs) and background (the remaining solar features and all off-disk areas). In the first step, one or more enclosed shapes, collectively called the contour ($C$), are defined. This contour acts as an initial guess, with regions inside the contour ($C^{+}$) serving as the foreground and regions outside of the contour ($C^{-}$) serving as the background. In the second step, the contour is refined by evaluating every pixel along the boundary of the contour to determine if it is more similar to the pixels within the foreground or the background and redrawing $C$ accordingly. The second step is repeated until a stopping criterion, such as no evolution between iterations, is met.
In the standard formulation of ACWE, the process of determining whether a pixel along the contour boundary belongs to $C^{+}$ or $C^{-}$ is decided using the ACWE energy functional:
@@
where $I$ is the image. This equation summarizes three “forces” that act on $C$, giving each force a user-defined weight that determines the relative importance in defining the final contour. The contour is manipulated in order to minimize this energy functional. The first force,
@@
attempts to minimize the length of the contour $\ell(C)$, and uses the user-defined weight $\mu$ to determine its relative importance. The second force,
@@
attempts to create a “homogeneous” foreground, where homogeneous is understood to mean a narrow range of intensities. This is done by comparing every pixel of $I(x,y)$ in $C^{+}$, to the mean intensity of the foreground $m^{+}$. This force is subject to the user-defined weight $\lambda^{\rm{H}+}$. The final force,
@@
mirrors the behavior of the second force (Equation 3), this time trying to create a homogeneous background by comparing every pixel in $C^{-}$ to the mean intensity of the background $m^{-}$. It is subject to the user-defined weight $\lambda^{\rm{H}-}$.
In both Boucheron et al. (2016) and Grajeda et al. (2023), the initial contour (seed) is defined using an intensity threshold applied to a 193 Å EUV image that has been corrected for limb brightening using the method from Verbeeck et al. (2014). This threshold defines all on-disk pixels with an intensity $\leq\alpha\times m_{\rm QS}$ (where $\alpha$ is a user-defined parameter [typically $0.3$] and $m_{\rm QS}$ is an estimate of the mean intensity of the QS) as CHs. The contour is evolved using the same 193 Å EUV image, after masking off-disk areas from the algorithm, and is stopped when either no evolution has occurred between iterations or the only evolution that occurs consists of pixels along the boundary that oscillate between foreground and background.
@@ 2.2 Dataset and Metrics
For this work we will be utilizing the dataset we previously developed in Grajeda et al. (2023). This dataset contains AIA Level 1 EUV images at 94, 131, 171, 193, 211, 304, and 335 Å, and the temporally corresponding 720-second line-of-sight (LOS) HMI magnetograms, collected at a one-hour cadence from Carrington rotations (CRs) 2099, 2100, 2101, and 2133. We note that all magnetogram data have an observation time within ±$\pm 2$ s of the record time for the corresponding EUV data. We also note that LOS magnetograms may report inaccurate magnetic field at the solar limb, but that our method should be able to equivalently use radial magnetic field from vector magnetograms (an area of future research). Our workflow will be the reverse of the process in Grajeda et al. (2023), tuning our proposed variant of the ACWE algorithm to operate during CR 2133, which has a high level of filament activity, to ensure that this version of the algorithm can minimize false positive detections. Once the algorithm is tuned, the observations from CRs 2099-2101 will be used to verify that the algorithm continues to perform effectively during the periods of low solar activity where it previously excelled.
To allow for an accurate comparison between Grajeda et al. (2023) and this work, we will be utilizing the same four metrics: intersection over union (IOU), structural similarity index measure (SSIM), global consistency error (GCE), and local consistency error (LCE). Both IOU and SSIM report a value of 1 when two segmentations are identical, with IOU serving as a stringent method that reports the overlap of the two segmentations normalized by total area and SSIM providing a perceptual measure that is more consistent with the human visual system (Wang et al., 2004). Both GCE and LCE report a value of 0 when two segmentations are identical. GCE and LCE are designed to report minimal or no error when one segmentation is a refinement of the other, with GCE serving as a more stringent metric (Martin et al., 2001).
@@ 3 Quantifying Unipolarity via Active Contour Kinetics
Incorporation of magnetic field data into ACWE evolution requires two adjustments to the standard formulation of ACWE. First, the ACWE algorithm must be altered to allow for vector-valued images so that EUV and HMI magnetogram data can be simultaneously considered. Second, the ACWE energy functional must be altered to maximize unipolarity as expressed by the magnetic field data while still maximizing homogeneity in the EUV observation(s).
@@ 3.1 ACWE for Vector-Valued Images
This new formulation evaluates the homogeneity of $C^{+}$ and $C^{-}$ on a per-channel basis by comparing the pixels (foreground and background, respectively) in each channel $I_{i}$ of the vector-valued image $I$ to the channel-specific means $m^{+}_{i}$ and $m^{-}_{i}$. Within this formulation the weights for the two homogeneity parameters are also defined on a per-channel basis as $\lambda^{\rm{H}+}_{i}$ and $\lambda^{\rm{H}-}_{i}$ (foreground and background, respectively), allowing the user to adjust the relative influence each channel’s foreground and background has on the final contour. It should be noted that when the total number of channels $N=1$, the formulation in Equation 5 reduces to the original ACWE energy functional in Equation 1.
@@ 3.2 Incorporating Magnetic Unipolarity
As originally formulated, the vector-valued version of ACWE only considers the homogeneity of the underlying region within each channel. For this reason, the energy functional in Equation 5 must be modified to consider homogeneity in EUV only, while instead considering unipolarity in the corresponding magnetograms. Forcing ACWE to ignore homogeneity when evaluating the magnetic field can be accomplished by setting $\lambda^{\rm{H}+}_{j}=0$ and $\lambda^{\rm{H}-}_{j}=0$ where $j$ is the index of the HMI magnetogram data within the vector-valued image $I$. From there, a metric that minimizes when a region becomes more unipolar can be added to the vector-valued energy functional (Equation 5).
We use the metric of Ko et al. (2014) as the basis for our measure of unipolarity. Ko et al. (2014) define unipolarity as:
@@
[…]
In order to incorporate Equation 6 into the vector-valued energy functional in Equation 5, the effect of moving a pixel at the contour boundary into or out of the foreground (CH) must be quantified. This means that both $\langle|B|\rangle$ and $|\langle B\rangle|$ must be calculated on a per-pixel basis. This process can be simplified by relying on the fact that the mean of a vector $x$ is defined as
@@
If the mean $m_{x_{N-1}}$ of a subset of $x(i)$ for $i=[1,N-1]$ has already been calculated, then the mean of the full ensemble for $i=[1,N]$ can be expressed as
@@
Using this formulation for the mean, the effect of a pixel $I(x,y)$ along $C$ on the unipolarity of the foreground (in the $i^{th}$ channel) can be expressed as
@@
where $a^{+}_{i}$ is the absolute mean of the foreground (of the $i^{th}$ channel), and $n^{+}$ is the number of pixels in the foreground. Note that we are using $\lambda^{\rm{U}+}_{i}$ to refer to the user-defined weight. Likewise, the unipolarity of the background can now be expressed as
@@
Adding Equation 9 and Equation 10 to the vector-valued energy functional (Equation 5) yields our new energy functional:
@@
Here, note that there are foreground and background means for the EUV channel(s) and the HMI channel.

### v2

@@ Abstract
Coronal holes (CHs) are magnetically open regions that allow hot coronal plasma to escape from the Sun and form the high-speed solar wind. This wind can interact with Earth’s magnetic field. For this reason, developing an accurate understanding of CH regions is vital for understanding space weather and its effects on Earth. The process of identifying CH regions typically relies on extreme ultraviolet (EUV) imagery, leveraging the fact that CHs appear dark at these wavelengths. Accurate identification of CHs in EUV, however, can be difficult due to a variety of factors, including stray light from nearby regions, limb brightening, and the presence of filaments (which also appear dark, but are not sources of solar wind). In order to overcome these issues, this work incorporates photospheric magnetic field data into a classical EUV-based segmentation algorithm based on the active contours without edges (ACWE) segmentation method. In this work magnetic field data are incorporated directly into the segmentation process, serving both as a method for removing non-CH regions in advance, and as a method to constrain evolution of the segmented CH boundary. This reduces the presence of filaments while allowing the segmentation to include CH regions that may be difficult to identify due to inconsistent intensities.
footnotetext: This version of the article has been accepted for publication, after peer review, but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at https://doi.org/10.1007/s11207-025-02536-7
@@ 1 Introduction
Due to the lower density and temperature of CH regions, a byproduct of the open magnetic field lines, CHs can be detected in extreme ultraviolet (EUV) imagery, where they appear as dark regions (Munro and Withbroe, 1972). This has resulted in a large number of detection schemes which rely on EUV imagery, including those developed by Krista and Gallagher (2009); Verbeeck et al. (2014); Lowder et al. (2014); Caplan et al. (2016); Boucheron et al. (2016); Garton et al. (2018); Hamada et al. (2018); Grajeda et al. (2023). Each of these methods addresses difficulties with the segmentation process, including inconsistent intensities caused by stray light from nearby regions, limb brightening, and instrument effects (Verbeeck et al., 2014; Caplan et al., 2016). In addition to these difficulties, the presence of filaments, which also appear dark in EUV but are not sources of solar wind, can result in the misclassification of regions (Krista and Gallagher, 2009; Reiss et al., 2023).
@@
The open magnetic field lines present in CHs emerge from highly unipolar regions unlike magnetic fields in filaments which are bipolar in nature. This has resulted in several methods that validate CH regions after detection through the use of magnetic field data, removing non-CH regions as a post-hoc process (Scholl and Habbal, 2008; Krista and Gallagher, 2009; Lowder et al., 2014; Garton et al., 2018; Hamada et al., 2018; Landeros et al., 2025). Direct application of magnetic field data in the initial segmentation of CHs remains limited, appearing predominantly in Jarolim et al. (2021) where it serves as an additional channel alongside seven EUV channels input to a network based on the U-Net architecture of Ronneberger et al. (2015). The work presented here proposes incorporating magnetic field data into the classical segmentation method first employed by Boucheron et al. (2016) and extended in Grajeda et al. (2023), which is based on Active Contours Without Edges (ACWE) (Chan and Vese, 2001). This ACWE-based method has been shown to effectively overcome inconsistent intensities present in CH observations for the majority of cases and is robust across short timeframes where CH evolution is expected to be minimal (Grajeda et al., 2023). However, like all other EUV-based methods, it can misidentify filament regions (Grajeda et al., 2023; Reiss et al., 2023). Through the addition of magnetic field data in both seeding and iterating of the ACWE process, the method proposed here reduces filament contamination, first by eliminating many filaments outright and second by reducing the area of filaments that cannot be outright removed. Additionally, incorporation of magnetic field better constrains the evolution to avoid including too much quiet Sun (QS) as CH.
@@
This paper is organized as follows: Section 2 provides an introduction to ACWE segmentation, outlines how it was adapted for CH detection, and introduces the datasets. Section 3 outlines the incorporation of magnetic field data into ACWE evolution and explores the resulting effects. Section 4 outlines the incorporation of magnetic field data into the initial seeding to help remove filament contamination outright. The final pipeline is discussed and validated in Section 5. Conclusions and discussion of future work are provided in Section 6.
@@ 2.1 Active Contours Without Edges (ACWE)
The detection method proposed here is an extension of the ACWE algorithm developed by Chan and Vese (2001) and first applied to CH detection in Boucheron et al. (2016). ACWE is a two-step process that separates an image into foreground (CHs in our case) and background (remaining solar features and off-disk areas). In the first step, one or more enclosed shapes, collectively called the contour ($C$), are defined. This contour acts as an initial guess, with regions inside the contour ($C^{+}$) serving as the foreground and regions outside the contour ($C^{-}$) as the background. In the second step, the contour is refined by evaluating every pixel along the contour boundary to determine if it is more similar to the pixels within $C^{+}$ or $C^{-}$ and redrawing $C$ accordingly. The second step is repeated until a stopping criterion is met.
In the standard formulation of ACWE, determining whether a pixel along the contour boundary belongs to $C^{+}$ or $C^{-}$ is decided using the energy functional:
@@
where $I$ is the image. This equation summarizes three “forces” that act on $C$, giving each force a user-defined weight that determines its relative importance. The contour is manipulated to minimize the functional. The first force,
@@
attempts to minimize the length of the contour $\ell(C)$, and has user-defined weight $\mu$ to determine its relative importance. The second force,
@@
attempts to create a “homogeneous” foreground, where homogeneous is understood to mean a narrow range of intensities. This is done by comparing every pixel of $I(x,y)$ in $C^{+}$, to the mean intensity of the foreground $m^{+}$. This force is subject to user-defined weight $\lambda^{\rm{H}+}$. The final force,
@@
mirrors the behavior of the second force (Equation 3), this time trying to create a homogeneous background by comparing every pixel in $C^{-}$ to the mean intensity of the background $m^{-}$. It is subject to user-defined weight $\lambda^{\rm{H}-}$. By varying the relationship between the two homogeneity parameters $\lambda^{\rm{H}+}$ and $\lambda^{\rm{H}-}$, the user can prioritize enforcing homogeneity in either $C^{+}$ or $C^{-}$. It should be noted that the foreground, which consists of CHs, is expected to be more homogeneous than the aggregate of all remaining solar features (including QS, filaments, and active regions). For this reason, both Boucheron et al. (2016) and Grajeda et al. (2023) enforce homogeneity in $C^{+}$, and allow for inhomogeneity in $C^{-}$, by co-defining $\lambda^{\rm{H}+}$ and $\lambda^{\rm{H}-}$ as a ratio $\lambda^{\rm{H}+}/\lambda^{\rm{H}-}\geq 10$.
In both Boucheron et al. (2016) and Grajeda et al. (2023), the initial contour (seed) is defined using an intensity threshold applied to a 193 Å EUV image that has been corrected for limb brightening using the method from Verbeeck et al. (2014). This threshold defines all on-disk pixels with an intensity $\leq\alpha\times m_{\rm QS}$ (where $\alpha$ is a user-defined parameter [typically $0.3$] and $m_{\rm QS}$ is an estimate of the mean intensity of the QS) as CHs. The contour is evolved using the same image, after masking off-disk areas from the algorithm. The process is stopped if no evolution occurs between iterations or if the only evolution that occurs consists of pixels along the boundary that oscillate between $C^{+}$ and $C^{-}$.
@@ 2.2 Dataset and Metrics
2.2.1 Primary Dataset and Metrics
For this work we utilize the dataset we previously developed in Grajeda et al. (2023). This dataset contains AIA Level 1 EUV images at 94, 131, 171, 193, 211, 304, and 335 Å, and the temporally corresponding 720-second line-of-sight (LOS) HMI magnetograms, collected at a one-hour cadence from Carrington rotations (CRs) 2099, 2100, 2101, and 2133. All magnetogram data have an observation time within $\pm 2$ s of the record time for the corresponding EUV data. We note that LOS magnetograms may report inaccurate magnetic field at the solar limb, but our method should be able to equivalently use radial magnetic field from vector magnetograms (an area of future research). Our workflow will be the reverse of the process in Grajeda et al. (2023), tuning our proposed variant of ACWE on CR 2133 which has a high level of filament activity, to ensure that this algorithm can minimize false positive detections. Then, observations from CRs 2099-2101 will be used to verify that the algorithm continues to perform effectively during periods of low solar activity where it previously excelled.
To accurately compare between Grajeda et al. (2023) and this work, we utilize the same four metrics: intersection over union (IOU), structural similarity index measure (SSIM), global consistency error (GCE), and local consistency error (LCE). Both IOU and SSIM report a value of 1 when two segmentations are identical, with IOU serving as a stringent metric quantifying overlap between the two segmentations normalized by total area and SSIM providing a perceptual measure more consistent with human vision (Wang et al., 2004). Both GCE and LCE report a value of 0 when two segmentations are identical. GCE and LCE report minimal or no error when one segmentation is a refinement of the other, with GCE being a more stringent metric (Martin et al., 2001).
2.2.2 Secondary Datasets and Metrics
In addition to the dataMSEE and MSEE and set of Grajeda et al. (2023), used to develop this method, two other datasets will be used in validation. The first dataset is from Reiss et al. (2023). This dataset provides labels for both CHs and filaments over 29 observations from July 2014 through June 2019. This allows for a direct comparison between the previous formulation of ACWE (Boucheron et al., 2016; Grajeda et al., 2023) and the formulation presented here. For this comparison we provide the total number of CHs, filaments, and other objects detected by each method. We also provide a comparison of the total area, in pixels and Mm2, for each filament still present in the updated formulation.
Second, we have collected a dataset with daily cadence that contains observations from the start of CR 2099 through the end of CR 2294 (13 July 2010 through 2 March 2025). This dataset is used to verify behavior of the proposed method by measuring the total area in Mm2 and magnetic field underlying each segmented CH. This determines if the behavior of objects detected by our proposed method is consistent with expectations over the portions of solar cycles 24 and 25 that SDO has captured. For this dataset we collect the seven AIA observations with record time closest to 12:00:00 on each day for which all observations have QUALITY key of 0 and there exists a 720-second LOS HMI magnetogram with QUALITY key of 0 and observation time within $\pm 2$ s of the AIA record time. We also collect that 720-second LOS HMI magnetogram.
@@ 3 Quantifying Unipolarity via Active Contour Kinetics
Incorporation of magnetic field data into ACWE evolution requires two adjustments to the standard formulation of ACWE. First, the ACWE algorithm must be altered to allow for vector-valued images so that EUV and magnetogram data can be simultaneously considered. Second, the ACWE energy functional must be altered to maximize unipolarity as expressed by the magnetic field data while still maximizing homogeneity in the EUV observation(s).
@@ 3.1 ACWE for Vector-Valued Images
This new formulation evaluates the homogeneity of $C^{+}$ and $C^{-}$ on a per-channel basis by comparing the pixels (foreground and background, respectively) in each channel $I_{i}$ of the vector-valued image $I$ to the channel-specific means $m^{+}_{i}$ and $m^{-}_{i}$. Within this formulation the weights for the two homogeneity parameters are also defined on a per-channel basis as $\lambda^{\rm{H}+}_{i}$ and $\lambda^{\rm{H}-}_{i}$ (foreground and background, respectively), allowing the user to adjust the relative influence of each channel’s foreground and background. It should be noted that when the total number of channels $N=1$, Equation 5 reduces to Equation 1.
@@ 3.2 Incorporating Magnetic Unipolarity
The vector-valued version of ACWE only considers homogeneity within each channel. For this reason Equation 5 must be modified to consider homogeneity in EUV only, while considering unipolarity in the corresponding magnetograms. Forcing ACWE to ignore homogeneity when evaluating the magnetic field can be accomplished by setting $\lambda^{\rm{H}+}_{j}=0$ and $\lambda^{\rm{H}-}_{j}=0$ where $j$ is the index of the magnetogram data within the vector-valued image $I$. From there, a metric that minimizes when a region becomes more unipolar can be added to Equation 5.
We use the unipolarity metric of Ko et al. (2014):
@@
where $\langle\cdot\rangle$ denotes mean, $|\cdot|$ denotes absolute value, and $B$ is the magnetic field. This metric is bound to the range $[0,1]$, where 0 indicates a purely unipolar region and 11 indicates a purely bipolar region. This ensures that increasing unipolarity will minimize the energy functional.
In order to incorporate Equation 6 into Equation 5, the effect of moving a pixel at the contour boundary into or out of the foreground (CH) must be quantified. This means that both $\langle|B|\rangle$ and $|\langle B\rangle|$ must be calculated on a per-pixel basis. This can be achieved by relying on the fact that the mean of a vector $x$ is
@@
If the mean $m_{x_{N-1}}$ of a subset of $x(i)$ for $i=1,\ldots,N-1$ is already calculated, then the mean of the full ensemble for $i=1,\ldots,N$ is
@@
Using Equation 8, the effect of a pixel $I(x,y)$ along $C$ on the unipolarity of the foreground (in the $i^{th}$ channel) can be expressed as
@@
where $a^{+}_{i}$ is the absolute mean of $C^{+}$ (of the $i^{th}$ channel), and $n^{+}$ is the number of pixels in $C^{+}$. Note that we are using $\lambda^{\rm{U}+}_{i}$ to refer to the user-defined weight. Likewise, the unipolarity of the background can be expressed as
@@
Adding Equations 9 and 10 to Equation 5 yields our new energy functional:
@@
In this formulation, we mirror the structure of the homogeneity-only version of ACWE for a vector-valued image (Equation 5), providing unipolarity constraints for both the foreground and background. This allows the user to prioritize either a unipolar foreground or background in a given channel ($i$) by varying the relationship between $\lambda^{\rm{U}+}_{i}$ and $\lambda^{\rm{U}-}_{i}$. Since we wish to enforce unipolarity in the foreground (CH region), and not in the background, the relationship $\lambda^{\rm{U}+}_{j}/\lambda^{\rm{U}-}_{j}$ should be >1>1. We refer to the formulation in Equation 11 as “QUACK” (Quantifying Unipolarity via Active Contour Kinetics), and use “ACWE” for the methods in Boucheron et al. (2016); Grajeda et al. (2023).